# Denoising weak lensing mass maps with diffusion model: systematic comparison with generative adversarial network

### v1

@@ 4.4.1 Power spectrum and cross-correlation
where $\tilde{\kappa}$ denotes the convergence field in Fourier space. Since the convergence maps are normalised with the variance, we define the normalized power spectrum $\tilde{C}(\ell)$ as $\tilde{C}(\ell)=C(\ell)/\sigma^{2}$, where $\sigma^{2}$ is the variance of the convergence map. The power spectra are computed with 20 bins equally spaced in logarithmic space for the range $\ell\in[200,10^{4}]$.
@@
$\Delta\tilde{C}(\ell)/\mathrm{Var}[\tilde{C}(\ell)]^{1/2}\equiv\frac{\tilde{C}_{\mathrm{denoised}}(\ell)-\tilde{C}_{\mathrm{truth}}(\ell)}{\mathrm{Var}[\tilde{C}_{\mathrm{truth}}(\ell)]^{1/2}}.$ (33)
As described in Section 3.3, there are five samples from GAN and DM from the single noisy map, and the thin solid lines correspond to the mean power spectra of 1,000 generated maps for each of the five samples. The thick solid lines correspond to the average of all 5,000 maps. Overall, DM outperforms GAN with regard to power spectrum reconstruction; DM can reproduce the power spectrum within $0.1\sigma$ for $\ell\lesssim 4000$. Though the noise power spectrum dominates for $\ell\gtrsim 2000$, DM accurately reconstructs the power spectrum at such scales. On the other hand, the range where GAN reproduces the correct power spectrum is limited to the large scale ($\ell\lesssim 1000$). Furthermore, the variance among the five networks for GAN is quite large even at the intermediate scale ($\ell\sim 1000$). Averaging over five samples results in a reasonable estimate of the power spectrum but the deviation for some models reaches $0.5\sigma$. In contrast, the five samples of DM indicate less variance; all five samples perform well up to $\ell\simeq 4000$.
@@
[…]
As naively expected, both GAN and DM reconstruct the target fields at large scales to some extent, and the performance worsens for smaller scales. In terms of the cross-correlation coefficients, a significant difference between GAN and DM cannot be seen.
@@ 4.4.2 Bispectrum
Next, we focus on the non-Gaussian statistics, the bispectrum. The bispectrum is the counterpart in Fourier space of the three-point correlation function and conveys information beyond the two-point correlation functions (see, e.g., Takada & Jain 2004). The bispectrum of the convergence field $B(\bm{\ell}_{1},\bm{\ell}_{2},\bm{\ell}_{3})$ is defined as
@@
[…]
In this study, we employ the reduced bispectrum $\tilde{B}$ which are computed from the normalised convergence maps. The reduced bispectrum is computed based on the estimator of Scoccimarro (2015) with the multipole range $\ell\in[200,10000]$ with $20$ logarithmically-spaced bins. Specifically, we consider three configurations: equilateral ($\ell_{1}=\ell_{2}=\ell_{3}$), flatten ($\ell_{1}=\alpha\ell_{2}=\alpha\ell_{3}$ with $\alpha=0.556$), and squeezed ($\ell_{1}=\ell_{2}>\ell_{3}$ with $\ell_{3}=220.549$) configurations.
Figures 7, 8, and 9 show the bispectra for the equilateral, flattened, and squeezed configurations, respectively. DM reproduces the bispectrum for all configurations within $0.05\sigma$ level down to small scales $\ell\lesssim 5000$ while the accuracy of GAN exceeds $1\sigma$ at relatively large scales. In addition, variations across different samples of DM are clearer than in the power spectrum. For GAN, variations similar to the case in the power spectrum appear across different networks and the fluctuation amplitude is larger than that of DM. The features do not vary for different configurations.
@@ 4.4.3 Probability distribution function
Next, we focus on the one-point probability density function (PDF) of the denoised convergence maps. The PDF also contains information beyond the two-point statistics and can be easily measured in actual observations (Thiele et al. 2023). We measure the PDF of the normalised convergence map with the range $\mathrm{SNR}\in[-5,15]$ with 100 linearly spaced bins and the regions closer to the edges within $2\theta_{\mathrm{G}}$ are excluded. The PDFs of the denoised maps with GAN and DM are shown in Figure 10. The reconstruction accuracy of DM is at least $0.2\sigma$ for the entire range. All five samples show a similar trend and this consistency demonstrates the robustness of the sampling process inherent to the DM. On the other hand, the results of GAN yield larger variability among five trained networks. All GAN samples show deviations greater than $0.25\sigma$ in some range, with significant deviations around $\mathrm{SNR}\simeq 0$ and $5$.
@@ 4.4.4 Peak and minima counts
Here, we discuss the peak and minima counts. Peaks (minima) are defined as pixels with values larger (smaller) than their eight surrounding pixels, and thus, the identification of peaks and minima in the actual analysis is straightforward. High peaks have the clear physical origin; massive dark matter halos are the source of the high signal (Yang et al. 2011). The low peaks originate from the line-of-sight alignment of smaller halos. In contrast, minima are negative peaks and mostly correspond to the underdense regions. Therefore, peaks and minima are employed to search overdense and underdense structures in the Universe. Furthermore, the counts of peaks and minima as a function of the height or depth can be used to constrain cosmological parameters (Coulton et al. 2020; Marques et al. 2024).
In this study, the counts of peaks and minima are measured in the range $\mathrm{SNR}\in[-5,15]$ with 100 linearly spaced bins and the regions closer to the edges within $2\theta_{\mathrm{G}}$ are excluded. The results are presented in Figure 11. DM can reconstruct the peak counts for $\mathrm{SNR}\gtrsim 3$ with high accuracy ($\lesssim 0.1\sigma$) but fails at low peaks. As shown in previous results, the reconstruction at small scales is harder both for DM and GAN, and that leads to low accuracy for low peaks. GAN can also reconstruct peak counts at very high peaks ($\mathrm{SNR}\gtrsim 8$) but the performance is worse than DM for the moderately high peaks ($\mathrm{SNR}\gtrsim 4$). DM outperforms GAN for every statistics, but it turns out that DM is not effective in reconstructing the minima counts. The minima counts with higher depth ($\mathrm{SNR}\gtrsim 2$) are well reproduced but for those with lower depth ($\mathrm{SNR}\lesssim 1$), the accuracy degrades to $1\sigma$ level. Both for peak and minima counts, DM can reconstruct the counts for lower SNR and enables one to explore more information.
@@ 4.4.5 Scattering transform
Scattering transform, originally proposed by Mallat (2012), is a method for extracting information from high-dimensional data, and its application in cosmology has been gaining attention in recent years. The scattering transform coefficients can capture a significant amount of non-Gaussianity in convergence maps and have been shown to provide stronger constraints on cosmological parameters compared to the power spectrum, peak counts, and even convolutional neural network (CNN) approach (Cheng et al. 2020). The scattering transform applies wavelet convolutions and modulus operations to the input image, transforming it into new representations for statistical analysis. Due to its use of convolution and non-linear modulus operations, it shares a structural similarity with CNNs. While CNNs have recently been able to improve parameter constraints through advances in ML (Sharma et al. 2024), the scattering transform has the advantage of being non-trainable, making it a useful alternative.
In this study, we follow the approach presented in Cheng et al. (2020) and utilize scattering package  [^3] https://github.com/SihaoCheng/scattering_transform to compute the scattering transform coefficients. Given an input convergence map $I_{0}$, we apply the first-order scattering transform to obtain the first-order representation $I_{1}$:
@@
where $\star$ denotes the convolution operation, $\Psi^{j_{1},l_{1}}$ is the wavelet with scale index $j_{1}$ and orientation index $l_{1}$, and ||$\cdot$|| ⋅ | denotes the modulus of a complex field. As the wavelet, we employ Morlet wavelets, which are defined as
@@
Cheng et al. (2020) indicate that higher-order coefficients, i.e., third-order and beyond, contribute to little additional cosmological information. Therefore, we restrict our analysis to the first- and second-order scattering transform coefficients.
From the fields $I_{1}$ and $I_{2}$ obtained via the scattering transform, we extract statistical quantities. Following previous works (Cheng et al. 2020; Shirasaki & Ikeda 2024), we compute the spatial average of the field. Since the convergence map is isotropic, we further take an average over the directional index $l$ to compute the scattering transform coefficients $s_{1}(j_{1})$ and $s_{2}(j_{1},j_{2})$:
@@
The first-order and second-order scattering transform coefficients are shown in Figure 12. Examining the first-order coefficients, we find that the DM and some GAN networks exhibit deviations that remain within $0.5\sigma$ across all scales $j_{1}$. For the second-order coefficients, while the model accurately reproduces results at large scales $j_{1}$, deviations become more noticeable at smaller scales. In contrast, GAN exhibits significant deviations depending on the specific network, and unlike DM, no clear pattern is observed for the second-order coefficients.
@@ 4.5 Stress tests
To evaluate the robustness and flexibility of the denosing process based on GAN and DM, we conduct stress tests by denoising maps with characteristics different from the training data. These stress tests mimic the real observations, where the true properties of lensing maps, such as the source redshift, are not exactly determined. We consider three cases: (i) maps with noise levels scaled by factors of $1.1$ and $0.9$, (ii) maps with the convergence field multiplied by $1.1$ and $0.9$, and (iii) maps with source galaxies at higher ($z_{s}=1.2$) and lower ($z_{s}=0.9$) redshifts. The training is conducted only once with the fiducial data set, and the maps in the stress tests have distinct features apart from the training data set. For GAN, sampling is performed for five networks for each test data, which results in 5,000 denoised maps. For DM, the variation among five samples is small, as demonstrated in Figure 5, and thus, only one sample is employed for stress tests. To highlight the changes relative to the fiducial test data, we show the normalised difference:
@@
where $\tilde{C}_{\mathrm{ref}}(\ell)$ denotes the reference power spectrum of the ground-truth maps from modified data, and $\tilde{C}(\ell)$ corresponds to the power spectrum of denoised maps in the stress test. Differently from the previous results, the reference power spectrum differs according to each stress test.
@@ 5 Conclusions
Weak lensing (WL) is the unique probe into the large-scale structures of the Universe and has been the primary science target in ongoing and future imaging cosmological surveys. Through measurements of the distortion of the shapes of galaxies, the matter distribution can be mapped with WL. The dominant contamination of the WL signal is the shape noise; the intrinsic shape of galaxies dilutes the signal due to the finite number of source galaxies. In this study, we apply machine learning (ML) techniques to remove the shape noise from observed noisy WL convergence maps. This noise removal process is referred to as denoising. To this end, we employ image-to-image translation approaches: generative adversarial network (GAN; Goodfellow et al. 2014) and diffusion model (DM; Sohl-Dickstein et al. 2015). GAN has been the major approach as the generative model, and the previous study (Shirasaki et al. 2019) demonstrates that GAN is effective at denoising WL maps. In addition to GAN, we also investigate the denoising performance of DM, which outperforms GAN in various practical tasks such as inpainting, super-resolution, colourisation, etc. Furthermore, DM learns the probability distribution of denoised maps conditioned on the observed maps, and thus, the trained model can generate multiple images from a single noisy convergence map. This is another advantage of DM over GAN, which basically learns one-to-one mapping due to technical difficulty.
@@
In order to evaluate the denoising performance in a realistic situation, we make use of the large suite of mock WL maps $\kappa\mathrm{TNG}$ (Osato et al. 2021). We construct 40,000 pairs of noisy observed mass maps and noiseless true mass maps. Then, 39,000 maps are used to train model parameters in GAN and DM, and the remaining 1,000 maps are used as test data. DM can sample multiple realisations of denoised maps, and we sample five maps from each noisy map, which amounts to 5,000 denoised maps in total. Though GAN basically can only sample a single realisation, we train five networks of GAN from different initial weights, which mimics multiple sampling. As a result, there are also 5,000 denoised maps with GAN.
@@
First, we investigate the pixel-level correlation between denoised maps and corresponding ground-truth maps. Both GAN and DM can reproduce the true pixel value with statistical fluctuation to a similar extent. Next, we study how the peaks in mass maps are affected by denoising. In the denoised maps with GAN and DM, there are similar numbers of peaks in true maps. However, roughly $45\%$ of peaks with $\mathrm{SNR}>4$ in denoised maps are not associated with the peaks in the true map. Moreover, $40\%$ of peaks with $\mathrm{SNR}>4$ in true maps are mistakenly erased by denoising. Therefore, our results suggest that challenges remain in accurately identifying dark matter halos and galaxy clusters when using ML.
@@
Next, we address how cosmological statistics can be recovered by denoising. To begin with, we measure the angular power spectrum, which is the most fundamental statistics in WL analysis, of denoised maps with GAN and DM. Then, we measure the higher-order statistics: bispectrum, one-point PDF, peaks and minima counts, and scattering transform. These higher-order statistics contain information beyond the power spectrum and can be employed to place tighter constraints on cosmological parameters. We find that DM demonstrates better performance for the reconstruction of statistics with the only exception of low $\mathrm{SNR}$ minima counts. In addition, the statistics of denoised maps with DM exhibit smaller fluctuations among the five samples. On the other hand, GAN shows quite a large variance among the five networks, and thus, that leads to the caveat that even if the mean of five GAN networks looks successful in reproducing the statistics, the individual model may fail at reconstruction and the accuracy can degrade at $1\sigma$ level.
@@
Finally, we conduct stress tests to address the robustness and flexibility of the trained denoising models. We create additional three data sets: noise scaled by $0.9$ and $1.1$, convergence scaled by $0.9$ and $1.1$, and source redshifts shifted to $z_{s}=0.9$ and $z_{s}=1.2$. Then, we denoise these maps with the model trained with original training data. Hence, the training and test data sets are inconsistent. However, this situation happens in real observations, where noise properties and source galaxy redshifts are not necessarily precisely known. We measure the power spectrum for denoised maps in the three modified test data sets. Even if the training is conducted with different data set, denoising with GAN and DM functions to some extent, though accuracy worsens. Similarly to the fiducial data sets, the accuracy of DM is better than GAN, and the deviation due to the modified data set is also smaller, which suggests that DM is robust to the changes in the characteristics of the mass maps.
@@
In this work, we explore the capabilities of GAN and DM in the specific task of denoising WL mass maps. We reveal that in general, DM performs better than GAN, which is anticipated from the higher quality of generated images with DM in previous studies. Moreover, GAN is known to be subject to mode collapse in training but the training of DM is more stable. However, the computational cost of training and sampling is more demanding for DM, which hampers DM from generating a large number of samples. There is another challenge regarding the size of mass maps. The number of grids in mass maps is limited by computational requirements; more grids lead to a large number of weights and may result in longer training. In order to apply the denoising with ML-based approaches to the real observational data, which covers wider areas than the simulation data used in this study, some techniques such as splitting large and small scale modes are critical. We leave it for future work.

### v2

@@ 4.4.1 Power spectrum and cross-correlation
[…]
@@
$\Delta\tilde{C}(\ell)/\mathrm{Var}[\tilde{C}(\ell)]^{1/2}\equiv\frac{\tilde{C}_{\mathrm{denoised}}(\ell)-\tilde{C}_{\mathrm{truth}}(\ell)}{\mathrm{Var}[\tilde{C}_{\mathrm{truth}}(\ell)]^{1/2}},$ (34)
where $\mathrm{Var}[\tilde{C}_{\mathrm{truth}}(\ell)]$ is the variance of the normalised true power spectra over 1,000 maps. As described in Section 3.3, there are five samples from GAN and DM from the single noisy map, and the thin solid lines correspond to the mean power spectra of 1,000 generated maps for each of the five samples. The thick solid lines correspond to the average of all 5,000 maps. Overall, DM outperforms GAN with regard to power spectrum reconstruction; DM can reproduce the power spectrum within $0.1\sigma$ for $\ell\lesssim 6000$. Though the noise power spectrum dominates for $\ell\gtrsim 2000$, DM accurately reconstructs the power spectrum at such scales. On the other hand, the range where GAN reproduces the correct power spectrum is limited to the large scale ($\ell\lesssim 1000$). Furthermore, the variance among the five networks for GAN is quite large even at the intermediate scale ($\ell\sim 1000$). Averaging over five samples results in a reasonable estimate of the power spectrum but the deviation for some models reaches $0.5\sigma$. In contrast, the five samples of DM indicate less variance; all five samples perform well up to $\ell\simeq 6000$.
@@
$r(\ell)=\frac{\Re\left[\langle\tilde{\kappa}_{\mathrm{denoised}}(\ell)\tilde{\kappa}^{*}_{\mathrm{truth}}(\ell)\rangle\right]}{\sqrt{\langle|\tilde{\kappa}_{\mathrm{denoised}}|^{2}\rangle\langle|\tilde{\kappa}_{\mathrm{truth}}|^{2}\rangle}},$ (35)
where $\Re[\cdot]$ corresponds to the operation to take the real part and $\langle\cdot\rangle$ is the average over 1,000 maps. As naively expected, both GAN and DM reconstruct the target fields at large scales to some extent, and the performance worsens for smaller scales. In terms of the cross-correlation coefficients, a significant difference between GAN and DM cannot be seen.
@@ 4.4.2 Bispectrum
Next, we focus on the non-Gaussian statistics, the bispectrum. The bispectrum is the counterpart in Fourier space of the three-point correlation function and conveys information beyond the two-point correlation functions (see, e.g., Takada2004). The bispectrum of the convergence field $B(\bm{\ell}_{1},\bm{\ell}_{2},\bm{\ell}_{3})$ is defined as
@@
$B(\bm{\ell}_{1},\bm{\ell}_{2},\bm{\ell}_{3})=\int_{0}^{\chi_{\mathrm{H}}}d\chi\,\frac{W_{\kappa}^{3}(\chi)}{r^{4}(\chi)}B_{\delta}\left(\frac{\bm{\ell}_{1}}{r(\chi)},\frac{\bm{\ell}_{2}}{r(\chi)},\frac{\bm{\ell}_{3}}{r(\chi)}\right).$ (37)
In this study, we employ the reduced bispectrum $\tilde{B}$ which are computed from the normalised convergence maps. The reduced bispectrum is computed based on the estimator of Scoccimarro2015 with the multipole range $\ell\in[200,10000]$ with $20$ logarithmically-spaced bins. Specifically, we consider three configurations: equilateral ($\ell_{1}=\ell_{2}=\ell_{3}$), flatten ($\ell_{1}=\alpha\ell_{2}=\alpha\ell_{3}$ with $\alpha=0.556$), and squeezed ($\ell_{1}=\ell_{2}>\ell_{3}$ with $\ell_{3}=220.549$) configurations.
Figures 7, 8, and 9 show the bispectra for the equilateral, flattened, and squeezed configurations, respectively. DM reproduces the bispectrum for all configurations within $0.05\sigma$ level down to small scales $\ell\lesssim 4000$ while the accuracy of GAN exceeds $1\sigma$ at relatively large scales. In addition, variations across different samples of DM are clearer than in the power spectrum. For GAN, variations similar to the case in the power spectrum appear across different networks and the fluctuation amplitude is larger than that of DM. The features do not vary for different configurations.
@@ 4.4.3 Probability distribution function
Next, we focus on the one-point probability density function (PDF) of the denoised convergence maps. The PDF also contains information beyond the two-point statistics and can be easily measured in actual observations (Thiele2023). We measure the PDF of the normalised convergence map with the range $\mathrm{SNR}\in[-5,15]$ with 100 linearly spaced bins and the regions closer to the edges within $2\theta_{\mathrm{G}}$ are excluded. The PDFs of the denoised maps with GAN and DM are shown in Figure 10. The reconstruction accuracy of DM is at least $0.2\sigma$ for the entire range. All five samples show a similar trend and this consistency demonstrates the robustness of the sampling process inherent to the DM. On the other hand, the results of GAN yield larger variability among five trained networks. All GAN samples show deviations greater than $0.25\sigma$ in some range, with significant deviations around $\mathrm{SNR}\simeq 0$ and $5$.
@@ 4.4.4 Peak and minima counts
Here, we discuss the peak and minima counts. Peaks (minima) are defined as pixels with values larger (smaller) than their eight surrounding pixels, and thus, the identification of peaks and minima in the actual analysis is straightforward. High peaks have the clear physical origin; massive dark matter halos are the source of the high signal (Yang2011). The low peaks originate from the line-of-sight alignment of smaller halos. In contrast, minima are negative peaks and mostly correspond to the underdense regions. Therefore, peaks and minima are employed to search overdense and underdense structures in the Universe. Furthermore, the counts of peaks and minima as a function of the height or depth can be used to constrain cosmological parameters (Coulton2020; Marques2024).
In this study, the counts of peaks and minima are measured in the range $\mathrm{SNR}\in[-5,15]$ with 100 linearly spaced bins and the regions closer to the edges within $2\theta_{\mathrm{G}}$ are excluded. The results are presented in Figure 11. DM can reconstruct the peak counts over the entire $\mathrm{SNR}$ range with an accuracy of $0.1\sigma$. In contrast, GAN can reproduce the peak counts only at very high peaks ($\mathrm{SNR}\gtrsim 8$), and its overall performance is worse than DM. For the minima counts, DM can also reproduce them with an accuracy of $0.2\sigma$ across the entire $\mathrm{SNR}$ range. Each GAN sample shows deviations of roughly $0.5\sigma$ at lower depth ($\mathrm{SNR}\lesssim 2$) and the mean still deviates by $0.5\sigma$ around $\mathrm{SNR}\simeq 0$. The reconstruction accuracy at low $\mathrm{SNR}$ strongly depends on the noise scheduling and we find that the quadratic scheduling, which is our fiducial choice, yields the best performance. The comparison between different noise scheduling schemes is presented in Appendix C.
@@ 4.4.5 Scattering transform
Scattering transform, originally proposed by Mallat2012, is a method for extracting information from high-dimensional data, and its application in cosmology has been gaining attention in recent years. The scattering transform coefficients can capture a significant amount of non-Gaussianity in convergence maps and have been shown to provide stronger constraints on cosmological parameters compared to the power spectrum, peak counts, and even convolutional neural network (CNN) approach (Cheng2020). The scattering transform applies wavelet convolutions and modulus operations to the input image, transforming it into new representations for statistical analysis. Due to its use of convolution and non-linear modulus operations, it shares a structural similarity with CNNs. While CNNs have recently been able to improve parameter constraints through advances in ML (Sharma2024), the scattering transform has the advantage of being non-trainable, making it a useful alternative.
In this study, we follow the approach presented in Cheng2020 and utilize scattering package [^5] https://github.com/SihaoCheng/scattering_transform to compute the scattering transform coefficients. Given an input convergence map $I_{0}$, we apply the first-order scattering transform to obtain the first-order representation $I_{1}$:
@@
where $\star$ denotes the convolution operation, $\Psi^{j_{1},l_{1}}$ is the wavelet with scale index $j_{1}$ and orientation index $l_{1}$, and $|\cdot|$ denotes the modulus of a complex field. As the wavelet, we employ Morlet wavelets, which are defined as
@@
Cheng2020 indicate that higher-order coefficients, i.e., third-order and beyond, contribute to little additional cosmological information. Therefore, we restrict our analysis to the first- and second-order scattering transform coefficients.
From the fields $I_{1}$ and $I_{2}$ obtained via the scattering transform, we extract statistical quantities. Following previous works (Cheng2020; Shirasaki2024), we compute the spatial average of the field. Since the convergence map is isotropic, we further take an average over the directional index $l$ to compute the scattering transform coefficients $s_{1}(j_{1})$ and $s_{2}(j_{1},j_{2})$:
@@
The first-order and second-order scattering transform coefficients are shown in Figure 12. Examining the first-order coefficients, we find that some GAN networks exhibit deviations that remain within $0.5\sigma$ across all scales $j_{1}$, whereas all DM samples show lower deviations, i.e., below $0.1\sigma$. The second-order coefficients show similar behaviour to the first-order ones: DM accurately reproduces results across all scales, and GAN exhibits significant deviations depending on the specific network.
@@ 4.5 Stress tests
To evaluate the robustness and flexibility of the denoising process based on GAN and DM, we conduct stress tests by denoising maps with characteristics different from the training data. These stress tests mimic the real observations, where the true properties of lensing maps, such as the source redshift, are not exactly determined. We consider three cases: (i) maps with noise levels scaled by factors of $1.1$ and $0.9$, (ii) maps with the convergence field multiplied by $1.1$ and $0.9$, and (iii) maps with source galaxies at higher ($z_{s}=1.2$) and lower ($z_{s}=0.9$) redshifts. These stress tests virtually correspond to applying denoising to WL surveys with different noise characteristics (case i) or source redshift distribution (case iii), or varying cosmological parameters (case ii). The training is conducted only once with the fiducial data set, and the maps in the stress tests have distinct features apart from the training data set. For GAN, sampling is performed for five networks for each test data, which results in 5,000 denoised maps. For DM, the variation among five samples is small, as demonstrated in Figure 5, and thus, only one sample is employed for stress tests. To highlight the changes relative to the fiducial test data, we show the normalised difference:
@@
[…]
@@ 5 Conclusions
Weak lensing (WL) is a unique probe into the large-scale structures of the Universe and has been the primary science target in ongoing and future imaging cosmological surveys. Through measurements of the distortion of the shapes of galaxies, the matter distribution can be mapped with WL. The dominant contamination of the WL signal is the shape noise; the intrinsic shape of galaxies dilutes the signal due to the finite number of source galaxies. In this study, we apply machine learning (ML) techniques to remove the shape noise from observed noisy WL convergence maps. This noise removal process is referred to as denoising. To this end, we employ image-to-image translation approaches: generative adversarial network (GAN; Goodfellow2014) and diffusion model (DM; Sohl-Dickstein2015). GAN has been the major approach as the generative model, and the previous study (Shirasaki2019) demonstrates that GAN is effective at denoising WL maps. In addition to GAN, we also investigate the denoising performance of DM, which outperforms GAN in various practical tasks such as inpainting, super-resolution, colourisation, etc. Furthermore, DM learns the probability distribution of denoised maps conditioned on the observed maps, and thus, the trained model can generate multiple images from a single noisy convergence map. This is another advantage of DM over GAN, which basically learns one-to-one mapping due to technical difficulty.
@@
In order to evaluate the denoising performance in a realistic situation, we make use of the large suite of mock WL maps $\kappa\mathrm{TNG}$ (Osato2021). We construct 40,000 pairs of noisy observed mass maps and noiseless true mass maps. Then, 39,000 maps are used to train model parameters in GAN and DM, and the remaining 1,000 maps are used as test data. DM can sample multiple realisations of denoised maps, and we sample five maps from each noisy map, which amounts to 5,000 denoised maps in total. Though GAN basically can only sample a single realisation, we train five networks of GAN from different initial weights, which mimics multiple sampling. As a result, there are also 5,000 denoised maps with GAN.
@@
First, we investigate the pixel-level correlation between denoised maps and corresponding ground-truth maps. Both GAN and DM can reproduce the true pixel value with statistical fluctuation to a similar extent. Next, we study how the peaks in mass maps are affected by denoising. In the denoised maps with GAN and DM, there are similar numbers of peaks in true maps. However, roughly $45\%$ of peaks with $\mathrm{SNR}>4$ in denoised maps are not associated with the peaks in the true map. Moreover, $40\%$ of peaks with $\mathrm{SNR}>4$ in true maps are mistakenly erased by denoising. Therefore, our results suggest that challenges remain in accurately identifying dark matter halos and galaxy clusters when using ML-based denoising.
@@
Next, we address how cosmological statistics can be recovered by denoising. To begin with, we measure the angular power spectrum, which is the most fundamental statistics in WL analysis, of denoised maps with GAN and DM. Then, we measure the higher-order statistics: bispectrum, one-point PDF, peaks and minima counts, and scattering transform. These higher-order statistics contain information beyond the power spectrum and can be employed to place tighter constraints on cosmological parameters. We find that DM demonstrates better performance for the reconstruction of all statistics. In addition, the statistics of denoised maps with DM exhibit smaller fluctuations among the five samples. On the other hand, GAN shows quite a large variance among the five networks, and thus, that leads to the caveat that even if the mean of five GAN networks looks successful in reproducing the statistics, the individual model may fail at reconstruction and the accuracy can degrade at $1\sigma$ level.
@@
Finally, we conduct stress tests to address the robustness and flexibility of the trained denoising models. We create additional three data sets: noise scaled by $0.9$ and $1.1$, convergence scaled by $0.9$ and $1.1$, and source redshifts shifted to $z_{s}=0.9$ and $z_{s}=1.2$. Then, we denoise these maps with the model trained with original training data. Hence, the training and test data sets are inconsistent. However, this situation happens in real observations, where noise properties and source galaxy redshifts are not necessarily precisely known. We measure the power spectrum for denoised maps in the three modified test data sets. Even if the training is conducted with different data set, denoising with GAN and DM functions to some extent, though accuracy worsens.
@@
In this work, we explore the capabilities of GAN and DM in the specific task of denoising WL mass maps. We reveal that in general, DM performs better than GAN, which is anticipated from the higher quality of generated images with DM in previous studies. Moreover, GAN is known to be subject to mode collapse in training but the training of DM is more stable. However, the computational cost of training and sampling is more demanding for DM, which hampers DM from generating a large number of samples. There is another challenge regarding the size of mass maps. The number of grids in mass maps is limited by computational requirements; more grids lead to a large number of weights and may result in longer training. In order to apply the denoising with ML-based approaches to the real observational data, which covers wider areas than the simulation data used in this study, some techniques such as splitting large- and small-scale modes are critical. Furthermore, several survey realisms are simplified in the presented work, e.g., non-uniform shape noise, survey masks, etc. These systematics should be taken into account to assess the capabilities of ML-based denoising approaches in realistic situations. We leave it for future work.
{ack}
This research was conducted using the FUJITSU Supercomputer PRIMEHPC FX1000 and FUJITSU Server PRIMERGY GX2570 (Wisteria/BDEC-01) at the Information Technology Center, The University of Tokyo.
Funding
This work was supported in part by JSPS KAKENHI Grant Number JP24H00215, JP25H01513, JP25K17380, JP25H00662 (K.O.) and JP24H00221 (M.S.).
Data availability
The data underlying this article will be shared on reasonable request to the corresponding author.